Circular No. 6286

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

TRANSIENT X-RAY BURSTER GRO J1744–28 C. Kouveliotou, Universities Space Research Association; J. Kom-

mers and W. H. G. Lewin, Massachusetts Institute of Technology; J. van Paradijs, University of Alabama, Huntsville (UAH) and University of Amsterdam; G. J. Fishman, Marshall Space Flight Center (MSFC), NASA; and M. S. Briggs, UAH; report: “Strong pulsations (∼ 50 percent modulation) near 2.1 Hz have been observed in the bursts from the burst source near the galactic center (IAUC 6275, 6276). Thus this unusual burst source and the transient pulsar GRO J1744–28 (IAUC 6284, 6285) are the same source. The hard x-ray bursts are therefore very likely the result of accretion instabilities on a strongly magnetized neutron star in a binary system.”

K. Hurley, Space Sciences Laboratory, University of California, Berkeley, on behalf of the Ulysses Gamma-Ray Burst Team; Kouveliotou, A. Harmon, Fishman, and Briggs, on behalf of the Gamma Ray Observatory BATSE team; van Paradijs, Kommers, and Lewin, write: “A refined position for the new repeating x-ray transient source (IAUC 6272) has been obtained using Ulysses and BATSE (specifically, one Ulysses/BATSE triggered event on Jan. 8). The position is an annulus whose halfwidth is 0^{o}.047 centered at α = 3^{h}02^{m}57^{s}, δ = −61^{o}34′07′′ (equinox 2000.0), with central radius 83^{o}.974. The intersection of this annulus with the positions reported on IAUC 6276 and 6284 defines a location that may be approximated by a box with the following coordinates: α = 17^{h}43^{m}22^{s}, δ = −28^{o}46′.7; 17^{h}45^{m}07^{s}, −28^{o}39′.2; 17^{h}45^{m}35^{s}, −28^{o}43′.1; 17^{h}43^{m}49^{s}, −28^{o}51′.0. This error box is smaller than, but nested within, the previous network synthesis localizations.”

M. H. Finger, Compton Observatory Science Support Center; R. B. Wilson, MSFC; and van Paradijs, communicate for the Compton Observatory BATSE team: “The pulse timing of the 467-ms-period x-ray pulsar GRO J1744–28 shows a binary orbital motion Doppler signature. A fit of frequencies observed between 1995 Dec. 21 and 1996 Jan. 10 yields the following parameters: orbital period 11.76(3) days; epoch of superior conjunction JD 2450079.71(2); asini = 2.59(3) light seconds; spin frequency 2.1410016(3) Hz (at 1995 Dec. 31.0 UT); and frequency derivative 8.6(5) × 10^{−12} Hz/s. We find a 90-percent confidence upper limit of 0.026 to the eccentricity. The x-ray mass function is f(M) = 1.31(4) × 10^{−4} M_{⊙}. For a 1.4-M_{⊙} neutron star and an inclination > 30^{o}, the companion mass is 0.065–0.14 M_{⊙}, with a Roche lobe radius of 4.1–5.2 solar radii.”

1996 January 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green